Does a reaction occur when aqueous solutions of lead(II) acetate and ammonium bromide are combined? yesno If a reaction does occur, write the net ionic equation. Use the solubility rules provided in the OWL Preparation Page to determine the solubility of compounds. Use the pull-down boxes to specify states such as (aq) or (s). If a box is not needed leave it blank. +
Answer #1

Pb(CH3COO)2(aq) + 2NH4Br(aq) = PbBr2(s) + 2NH4CH3COO(aq)

Pb2+(aq) + 2CH3COO-(aq) + 2NH4+(aq) + 2Br-(aq) = PbBr2(s) + 2NH4+(aq) + 2CH3COO-(aq) [ ionic equation ]

net ionic equation

Pb2+(aq) + 2Br-(aq) = PbBr2(s)

any query please comment

if satisfied please rate it thanks